Circular No. 7762

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVAE 2001ge, 2001gf, 2001gg, 2001gh, 2001gi, 2001gj G. Altavilla, S. Benetti, E. Cappellaro, A. Pastorello, M. Riello, M.

Turatto and L. Zampieri, Osservatorio di Padova; A. Clocchiatti and C. Contreras, Universidad Catolica de Chile; I. J. Danziger and P. Mazzali, Osservatorio di Trieste; and F. Patat, European Southern Observatory (ESO), report the discovery of six new supernovae on CCD frames taken on Nov. 11 (and confirmed on Nov. 18) during a search performed with the ESO/MPI 2.2-m telescope (+ Wide Field Imager). The supernovae were identified by comparing Nov. images with reference images taken (using the same equipment) on previous observing runs, and none were visible on images taken prior to last June (limiting mag 23.5–24.5). The V magnitudes in the table below are listed for Nov. 11 (V_{1}) and Nov. 18 (V_{2}).

SN  α_{2000}  δ_{2000}  V_{1}  V_{2}  z  Type
2001ge 20^{h}31^{m}14.^{s}44 −40^{o}35′16′′.0 22.5 21.4 0.22 Ia
2001gf 22 04 21.27 −18 21 56.8 20.7 20.0 0.20 Ia
2001gg 0 45 20.83 −29 45 12.2 23.7 23.4 0.61 II
2001gh 0 57 03.63 −27 42 32.9 21.6 20.9 0.16: II
2001gi 4 28 07.06 −36 21 45.2 21.0 21.7 0.20 Ia
2001gj 5 38 07.90 −23 42 34.4 22.4 22.8 0.27 II

SN 2001gf is located 2′′.07 west and 0′′.45 north of the center of its host galaxy; the other supernovae are within 1′′ of their host galaxy’s nucleus. Spectra obtained on Nov. 21 with the ESO 8.2-m MELIPAL Very Large Telescope (+ FORS1) reveal the types and redshifts listed above; from the spectral features, it is estimated that SNe 2001ge and 2001gf were a few days before maximum and SN 2001gi was two weeks after maximum. The spectrum of SN 2001gh shows a very blue, almost featureless continuum, consistent with a type-II supernova at very early phases.

COMET C/2000 WM_{1} (LINEAR) J. Watanabe, National Astronomical Observatory (NAO) of Japan,

reports the following antitail lengths and position angles for this comet from I-band CCD images taken with the NAO 0.50-m f/12 reflector (noting the earth’s passage through the orbital plane of the comet on Nov. 20.18 UT): Nov. 16.503, > 8′.5, 283^{o}; 17.542, > 6′.5, 300^{o}; 18.526, > 7′.9, 302^{o}; 19.555, > 6′.2, 317^{o}; 22.372, > 4′.7, 1^{o}. Small scale jet-like structure near the nucleus was also recognized on these images.

2001 November 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green